OpenStudy (aaronandyson):

log_{sqrt 2}(x + 1) = 2

OpenStudy (aaronandyson):

\[\log_{\sqrt 2}(x + 1) = 2\]

OpenStudy (aaronandyson):

Where 2 is log 100 i guess.

OpenStudy (ribhu):

x+1 = sqrt2 ^2

OpenStudy (ribhu):

x+1= 2 x=1

OpenStudy (ribhu):

property of logarithms

OpenStudy (anonymous):

Property of Logs: \[\log_a (b) = c \implies a^c = b\] Use it to solve for \(x\)..